How much work does the electric field do in moving a proton from a point with a potential of +170 V to a point where it is -70 V ? Express your answer in joules.Express your answer in electron volts.
Answer #1

workdone = change in potential energy

W = q*dV = 1.602*10^-19*240 = 384.48*10^-19 J

W = 384.48*10^-19*6.242*10^18 = 239.92 electron volt